OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x).

OpenStudy (anonymous):

can you solve \[7x-13=5\]?

OpenStudy (anonymous):

18/7

OpenStudy (anonymous):

steps are add 13 divide by 7

OpenStudy (anonymous):

so your inverse is \[f^{-1}(x)=\frac{x+13}{7}\]

OpenStudy (anonymous):

if you want a step by step procedure you already pretty much did it

OpenStudy (anonymous):

Okay that's what I got the first time

OpenStudy (anonymous):

start with \[7y-13=x\] solve for \(y\)

OpenStudy (anonymous):

\[7y-13=x\\ 7y=x+13\\ y=\frac{x+13}{7}\]